Village Finance Co. advanced three loans to Kamiko—$4,200 on June 20, $3,800 on August 25,and $3,300 on October 29. Simple interest at 8.25% was charged on all three loans, and all were repaid on December 31 when some bonds that she owned matured. What total amount was required to pay off the loans?


1
Expert's answer
2021-09-20T16:41:02-0400

Loan 1 = 4200$ taken on June 20


Loan 2 = 3800$ taken on August 25


Loan 3 = 3300$ taken on October 29


Loans will be repaid on 31st December


So the interest will be charged till 30th December


For loan 1 , the tenure of loan will be = June (11 days) including the 20th of June + July (31 days ) + aug (31days) + sep (30days) + Oct(31days) + Nov(30days) + dec(30days)


= 194 days


So time =t= 194/365


Interest charged =r= 8.25% annually , that is for 365 days


Interest charged on loan for its tenure = P×r×t /100 =


(4200×8.25×194)/(365×100) = 184.17


Amount required to close loan1 = 4200+184.17 = 4384.17 $


Similarly for loan2 = number of days = 7(aug) + 30(sep) + 31 (Oct) + 30 (Nov) + 30 (dec) = 128 days


Time = 128/365


Interest =( 3800×8.25×128)/(100×365) = 109.94$


Amount required to close loan2 = 3800+108.94= 3908.94


Similarly for loan3 tenure of loan = 3(Oct) + 30 (Nov) + 30 (dec) = 63 days


So interest charged = (3300×8.25×63)/(100×365) = 46.99


So amount required to close loan3 = 3300+46.99 = 3346.99


Total amount required to payoff the loans = 4384.17+ 3908.94 + 3346.99 = 11640.1 $